Below is a list of 30 marks: 74, 77, 35, 82, 91, 55, 56, 73, 72, 71 51, 99, 83, 78, 79, 62, 65, 56, 87, 88 39, 41, 93, 84, 75, 65, 67, 90, 80, 70 a) What percentile did someone who received a mark of 80 score in? b) What mark did someone receive who scored in the 73rd percentile?


1
Expert's answer
2021-10-04T18:25:07-0400

Sorting data in ascending order


35,39,41,51,55,56,56,62,65,65,67,70,71,72,73,74,75,77,78,79,80,82,83,84,87,88,90,91,93,99


(a)  i = (P/100) x N

Score of 80 has rank 21st so i = 20

20 = (P/100) x 30

P  = 67

So, it is 67th percentile


(b) i = (P/100) x N

  = 73/100 x 30

  = 21.9

  = 22

 

22nd data value is 82

So, 73rd percentile is 82